OpenStudy (anonymous):

I have been stuck on this question for like an hour z/(z^2-49) + 2/(z+7). Someone help please!!

OpenStudy (anonymous):

ok so what happens is this .... you take the fraction to the other side.. z/(z^2-49) = -2/(z+7) now you can factorize z^2-49 = (z+7)(z-7) so we get z/(z+7)(z-7) = -2/(z+7) now you multiply (z+7) on both sides so we get z/(z-7) = -2 this ends up being ... z = -2(z-7) z = -2z + 14 z+2z = 14 3z = 14 z = 14/3 so the final answer is z = 14/3

OpenStudy (anonymous):

ok, I think I got it!! Thank you so much!!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

There is no equation, therefore you do not get an answer. You are adding two fractions, so you must get a common denominator. Z^2 - 49 factors into (z - 7)(z + 7) The other denominator is (z+7) So both denominators need to be (z-7)(z+7) so you need to multiply the second fraction by (z-7)/(z-7) So you have (2(z-7))/(z+7)(z-7) which is (2z - 14)/(z + 7)(z - 7) Now you have two fractions with a common denominator so you are ready to add. z/((z + 7)(z - 7) + (2z - 14)/(z + 7)(z - 7) = (3z - 14)/((z + 7)(z - 7)